OpenStudy (anonymous):

Evaluate and write your answer in scientific notation:(4 x 102)(15 x 103) A. 60 x 106 B. 6 x 106 C. 6 x 105

OpenStudy (anonymous):

(4 x 10^2)(15 x 10^3)

OpenStudy (anonymous):

Just wanted to correct that

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 4 * 15 * 10^2 * 10^3 }\) \(\Large \color{MidnightBlue}{\Rightarrow 60 * 10 * 10 * 10 * 10 * 10 }\) \(\Large \color{MidnightBlue}{\Rightarrow 6000000 }\) \(\Large \color{MidnightBlue}{\Rightarrow 6 * 10^6 }\)

OpenStudy (cwrw238):

= 60 * 10^5 = 6 * 10^6

OpenStudy (anonymous):

(4 x 102)(15 x 103) = 60 x 10^5 = 6 x 10^6